Operation Summary for June 25, 2010 - June 28, 2010
Friday June 25
The day shift began with Operations monitoring the Tevatron (TeV) store 7922 (107.68E30) and the Recycler stash (348.3E10), with the Antiproton Source (Pbar) stacking (24.8E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTTB T-9777 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the day shift, Operations terminated store 7922 at 10:34 AM. Operators prepared D0 for a 1.5 hour access. Operators put the TeV through a dry squeeze and a wet squeeze. Operations reported that the MI-30 quadrupole power supply had tripped off three times; operators discovered a bad 24-volt PLC power supply. Operations reported that a MI RF station (MIRF6) bias power supply tripped a circuit breaker; RF technicians recovered the station. Operations put the TeV into shot setup.
During the evening shift, Operations established store 7925 at 5 PM with an average initial luminosity (AIL) of 348.8E30. Operations disabled the T-977 experiment at 10 PM.
Saturday June 26
The midnight shift began with Operations monitoring store 7925 (126.88E30) the stash (291.4E10), with Pbar stacking (15.1E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7925 at 4:03 AM, put the TeV into shot setup, and established store 7926 at 5:36 AM with an AIL of 345.9E30. Operations established beam to T-977 at 5:54 AM.
During the day shift, Operations reported that the Booster bunch length parameter (B:BL400M) was returning stale data; operators rebooted the node and the parameter resumed updating. At 10:50 AM, store 7926 aborted when the F0 utility yard breaker tripped off. There was no quench. The Duty Electrician reset the breaker and checked the current draw on the pumps and compressors. Operators prepared D0 for an access. Operations put the TeV into shot setup at 3:33 PM.
During the evening shift, Operations established store 7929 at 4:42 PM with an AIL of 236.55E30. Operations disabled beam to T-977 at 10:01 PM.
Sunday June 27
The midnight shift began with Operations monitoring store 7929 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-977 at 4 AM. At 4:07 AM, store 7929 aborted due to a TeV sector B11 separator spark. There was no quench. Operations established store 7930 at 6 AM with an AIL of 273.7E30. Operations reported short downtimes due to problems with LRF1 and 5, E:V108, and due to BRF9 and 10 vacuum problems.
During the day shift, a Switchyard F3 enclosure correction magnet (S:HT210) power supply tripped off on a ground fault indication and disabled the T-977 beam; it reset. The Booster East Anode Power Supply tripped off; it reset. Two Linac Klystron RF stations (KRF5 and 6) tripped off due to LCW flow faults; operators reset them.
During the evening shift, Operations disabled beam to T-977 at 10:03 PM. Operations established store 7931 at 11:53 PM with an AIL of 334.3E30.
Monday June 28
The midnight shift began with Operations monitoring store 7931 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations reported short downtimes due to KRF3, to a Booster RF ion pump (B:IPRF19), and to LRF2. Operations disabled the NuMI beam at 7:45 AM for a scheduled access.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 06/28/2010 email Fermilab|
|Security, Privacy, Legal|